# Envy-Free Allocation of Indivisible Goods via Noisy Queries

### v1

@@ C.2.4 Proof of Lemma 40 (Bound on $h$)
$=\frac{(c_{1}-c_{2})(e^{\prime}_{a}(c_{2})+e^{\prime}_{b}(c_{2}))}{(1+c_{1})(1+c_{2})}+\frac{1}{1+c_{1}}[e^{\prime}_{a}(c_{2})-e^{\prime}_{a}(c_{1})-c_{1}(e^{\prime}_{b}(c_{2})-e^{\prime}_{b}(c_{1}))]$ (92)
@@ C.4 The Case $q<m$
$\Delta<\min\{2\sigma^{2}m,\sqrt{2}\sigma m\}$ (101)
@@ C.4.2 Proof of Lemma 15 (Bound on $\overline{f}$)
We now observe that $\sum_{i\in S}|cu_{i}^{a}-u_{i}^{b}|$ is equal to the sum of values upon drawing $q$ items from a population of size $m$ without replacement, where the $i$-th value is $|cu_{i}^{a}-u_{i}^{b}|$. This allows us to apply Hoeffding’s inequality for sampling without replacement (e.g., see (Bardenet and Maillard,, 2015, Prop. 1.2)) to obtain
@@ C.4.3 Proof of Lemma 113 (Bound on $\overline{h}$)
Under the event $|v_{i}^{a}|\leq\sqrt{m}\log m$, we have for any consecutive $(c_{1},c_{2})$ (i.e, $c_{2}=c_{1}+\frac{1}{m^{3}}$) that
@@ D.1 Main Proof Steps via Auxiliary Results
and as a result, in order to establish a lower bound, it suffices to consider deterministic strategies. Note that this reduction to deterministic strategies is standard, in particular being a component of Yao’s minimax principle (e.g., see (Motwani and Raghavan,, 2010, Sec. 2.2.2)).
@@ D.3 Proof of Lemma 23 (Sufficient Conditions for Positive Envy)
Recall that $X_{[1:m]}$ are the binary variables defining the instance (see Table 1), $\widehat{X}_{[1:m]}$ are the corresponding “estimates” based on the allocation, and $\widehat{A}_{\varepsilon}$, $\widehat{B}_{\varepsilon}$, and $V_{\gamma}$ are defined in (157)–(159).
We define the following useful quantities:
@@
  • (178) follows since $N_{aa}+N_{ab}+N_{ba}+N_{bb}=\frac{m}{2}$ and $N_{ab}+N_{ba}=d_{H}(X_{[1:m/2]},\widehat{X}_{[1:m/2]})$.
@@ D.4 Proof of Lemma 24 (Lower Bound on Hamming Distance)
(Cai et al.,, 2023, Lemma 5) Under an i.i.d. prior on $Z_{1},\dotsc,Z_{m/2}$ and independent noise between queries, conditioned on any collection $\mathcal{D}$ of the $q$ query-outcome pairs, the variables $Z_{1},\dotsc,Z_{m/2}$ are independent.
@@ D.5 Proof of Lemma 25 (Lower Bound on $\Theta(\sqrt{m})$ Deviation Probability)
The proof of ˜24 concerned with the first half of the items (see Table 1), but the arguments up to (191) apply verbatim [^13] In this part of the analysis, the distinction between an item being “slightly favored by both agents” vs. “slightly favored by one and slightly disfavored by the other” is inconsequential. to the second half of the items upon replacing $\varepsilon$ by $\gamma$. Thus, adopting the shorthand $\zeta_{\gamma}(q,m)=\frac{\gamma\sqrt{2}}{\sigma}\sqrt{q\min\{q,\frac{m}{2}\}}$ for brevity, we have the following analog of (196):
@@
where we applied the Berry–Esseen theorem (Lemma 3 in Appendix A) to $\sum_{i\,:\,i>m/2}S_{i}X_{i}$ (with the substitutions $V_{T}\leftarrow\mathrm{Var}[\sum_{i>m/2}S_{i}X_{i}\,|\,\mathcal{D}]\geq C^{\prime}m$ due to (221), and $\Psi_{T}\leftarrow O(m)$ due to $S_{i}X_{i}\in[-1,1]$), and used the fact that the resulting Gaussian density is uniformly upper bounded by $\frac{1}{\sqrt{2\pi C^{\prime}m}}$. [^14] We note that $\frac{t}{\gamma}\cdot\frac{1}{\sqrt{2\pi C^{\prime}m}}$ serves as a uniform upper bound for the probability of the Gaussian random variable falling in any interval of length $\frac{t}{\gamma}$, so the precise value of ${\mathbb{E}}[\sum_{i>m/2}S_{i}X_{i}\,|\,\mathcal{D}]$ is not needed, and the shift by $\frac{c^{\prime}_{\mathcal{D}}}{2\gamma}$ in (222) is similarly inconsequential. Setting $t=\rho\gamma\sqrt{\pi C^{\prime}m}$, we find that (223) is at most $\frac{\rho}{\sqrt{2}}+O\big(\frac{1}{\sqrt{m}}\big)$, and is thus at most $\rho$ when $m$ is sufficiently large, thus implying the lemma with $K=\rho\sqrt{\pi C^{\prime}}$.

### v2

@@ C.2.4 Proof of Lemma 40 (Bound on $h$)
$=\frac{(c_{1}-c_{2})(e^{\prime}_{a}(c_{2})+e^{\prime}_{b}(c_{2}))}{(1+c_{1})(1+c_{2})}+\frac{1}{1+c_{1}}[e^{\prime}_{a}(c_{2})-e^{\prime}_{a}(c_{1})-c_{1}(e^{\prime}_{b}(c_{2})-e^{\prime}_{b}(c_{1}))].$ (92)
@@ C.4 The Case $q<m$
$\Delta<\min\{2\sigma^{2}m,\sqrt{2}\sigma m\}.$ (101)
@@ C.4.2 Proof of Lemma 15 (Bound on $\overline{f}$)
We now observe that $\sum_{i\in S}|cu_{i}^{a}-u_{i}^{b}|$ is equal to the sum of values upon drawing $q$ items from a population of size $m$ without replacement, where the $i$-th value is $|cu_{i}^{a}-u_{i}^{b}|$. This allows us to apply Hoeffding’s inequality for sampling without replacement (e.g., see ([)Prop. 1.2]bardenet2015concentration) to obtain
@@ C.4.3 Proof of Lemma 113 (Bound on $\overline{h}$)
Under the event $|v_{i}^{a}|\leq\sqrt{m}\log m$, we have for any consecutive $(c_{1},c_{2})$ (i.e., $c_{2}=c_{1}+\frac{1}{m^{3}}$) that
@@ D.1 Main Proof Steps via Auxiliary Results
and as a result, in order to establish a lower bound, it suffices to consider deterministic strategies. Note that this reduction to deterministic strategies is standard, in particular being a component of Yao’s minimax principle (e.g., see ([)Sec. 2.2.2]Mot10).
@@ D.3 Proof of Lemma 23 (Sufficient Conditions for Positive Envy)
Recall that $X_{[1:m]}$ are the binary variables defining the instance (see Table 1), $\widehat{X}_{[1:m]}$ are the corresponding “estimates” based on the allocation, and $\widehat{A}_{\varepsilon}$, $\widehat{B}_{\varepsilon}$, and $V_{\gamma}$ are defined in (157)–(159). We define the following useful quantities:
@@
  • (178) follows since $N_{aa}+N_{ab}+N_{ba}+N_{bb}=\frac{m}{2}$ and $N_{ab}+N_{ba}=d_{H}(X_{[1:m/2]},\widehat{X}_{[1:m/2]})$;
@@ D.4 Proof of Lemma 24 (Lower Bound on Hamming Distance)
([)Lemma 5]cai2023average Under an i.i.d. prior on $Z_{1},\dotsc,Z_{m/2}$ and independent noise between queries, conditioned on any collection $\mathcal{D}$ of the $q$ query-outcome pairs, the variables $Z_{1},\dotsc,Z_{m/2}$ are independent.
@@ D.5 Proof of Lemma 25 (Lower Bound on $\Theta(\sqrt{m})$ Deviation Probability)
The proof of ˜24 is concerned with the first half of the items (see Table 1), but the arguments up to (191) apply verbatim [^14] In this part of the analysis, the distinction between an item being “slightly favored by both agents” vs. “slightly favored by one and slightly disfavored by the other” is inconsequential. to the second half of the items upon replacing $\varepsilon$ by $\gamma$. Thus, adopting the shorthand $\zeta_{\gamma}(q,m)=\frac{\gamma\sqrt{2}}{\sigma}\sqrt{q\min\{q,\frac{m}{2}\}}$ for brevity, we have the following analog of (196):
@@
where we applied the Berry–Esseen theorem (Lemma 3 in Appendix A) to $\sum_{i\,:\,i>m/2}S_{i}X_{i}$ (with the substitutions $V_{T}\leftarrow\mathrm{Var}[\sum_{i>m/2}S_{i}X_{i}\,|\,\mathcal{D}]\geq C^{\prime}m$ due to (221), and $\Psi_{T}\leftarrow O(m)$ due to $S_{i}X_{i}\in[-1,1]$), and used the fact that the resulting Gaussian density is uniformly upper bounded by $\frac{1}{\sqrt{2\pi C^{\prime}m}}$. [^15] We note that $\frac{t}{\gamma}\cdot\frac{1}{\sqrt{2\pi C^{\prime}m}}$ serves as a uniform upper bound for the probability of the Gaussian random variable falling in any interval of length $\frac{t}{\gamma}$, so the precise value of ${\mathbb{E}}[\sum_{i>m/2}S_{i}X_{i}\,|\,\mathcal{D}]$ is not needed, and the shift by $\frac{c^{\prime}_{\mathcal{D}}}{2\gamma}$ in (222) is similarly inconsequential. Setting $t=\rho\gamma\sqrt{\pi C^{\prime}m}$, we find that (223) is at most $\frac{\rho}{\sqrt{2}}+O\big(\frac{1}{\sqrt{m}}\big)$, and is therefore at most $\rho$ when $m$ is sufficiently large, thus implying the lemma with $K=\rho\sqrt{\pi C^{\prime}}$.